40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Tue Jul 17 12:55:45 2018, Koji, Summary, General, Started pumping 6x
|
Tue Jul 17 18:16:38 2018, Steve, Summary, VAC, pumpdown 81 at 6 hrs 6x
|
Tue Jul 17 22:25:41 2018, Koji, Summary, VAC, pumpdown 81 at 6 hrs
|
Wed Jul 18 12:49:08 2018, Steve, Summary, VAC, pumpdown 81 at 6 +4.5hrs  
|
Wed Jul 18 17:36:50 2018, Steve, Summary, VAC, pumpdown 81 at 6 +9 hrs completed 
|
Thu Jul 19 01:56:25 2018, gautam, Summary, VAC, AUX pump shutdown
|
Thu Jul 19 11:01:03 2018, Steve, Summary, VAC, pd81 @ 2e-5 Torr
|
Thu Jul 19 13:35:30 2018, Steve, Summary, VAC, annuloses pumped
|
Mon Jul 23 09:58:52 2018, Steve, Summary, VAC, RGA scan at day 6
|
Mon Aug 6 09:34:02 2018, Steve, Update, VAC, RGA scan at day 20
|
Thu Aug 16 07:50:28 2018, Steve, Update, VAC, pumpdown 81 at day 30
|
Thu Aug 16 14:48:14 2018, Steve, Update, VAC, why do we need a root pump? 6x
|
Fri Nov 16 10:26:12 2018, Steve, Update, VAC, single viton O-rings
|
Fri Aug 24 10:51:23 2018, Steve, Update, VAC, pumpdown 81 at day 38
|
Thu Oct 4 08:25:50 2018, Steve, Update, VAC, rga scan pd81 at day 78
|
Fri Oct 12 08:27:05 2018, Steve, Update, VAC, drypump
|
Fri Oct 12 17:37:03 2018, Steve, Update, VAC, pressure gauge choices
|
Mon Oct 15 13:59:42 2018, Steve, Update, VAC, drypump replaced
|
Tue Nov 13 16:15:01 2018, Steve, Update, VAC, rga scan pd81 at day 119 
|
|
Message ID: 14077
Entry time: Tue Jul 17 12:55:45 2018
Reply to this: 14079
|
Author: |
Koji |
Type: |
Summary |
Category: |
General |
Subject: |
Started pumping |
|
|
[Steve, Koji, Gautam]
We started pumping down at ~12:15PM.
Vent finalization ~ YEND
- The table leveling was way off. This was adjusted by the balancing weight. (Attachment 1~3)
- The alignment of ETMY was not too much off. Just aligned it with the oplev spot on MEDM and this already made the green flashing.
- The Green TEM00 was maximized with ITMY and ETMY. This made the PSL IR flashing.
- The heater wires were checked. I found that one of the heater wires was touching the optical table via the cable shield. This is because the upper pins were shifted to the left side (Attachment 4&5). The pins were shifted and now all 4 cables are isolated form the table. I also checked the mutual resistance between the 4 terminals. They were measured to be isolated except two pairs that showed 4.4 Ohms and 4.0 Ohms (Attachment 6)
- The tools were removed from the chamber. The Y arm was still flashing.
- We closed the ETMY door.
Vent finalization ~ Vertex
- Found the ITMX stuck. Gautam came in and showed us his black magic to release the optic...
- This allowed us to align X arm. The green flash was found and the TEM00 flash was seen. This allowed us to see the PSL IR flash at the X end.
- PRM Refl was aligned. SRM was aligned with the oplev.
- The beam on the AS port was checked. The AS beam came out from the window.
- Closed the OMC chamber.
Pumping
- Started pumping with RP1 and RP3. (~12:15PM)
|
|
|
|
|
|
|
|
|